OpenStudy (anonymous):

HELP PLZ FAN AND MEDAL AND ANYTHING JUST HELP!!!! Choose the correct slope of the line that passes through the points (-4,8) and (-3,-6) Answers A. 7 B. -14 C. 0 D.14

OpenStudy (anonymous):

hi plz help...

OpenStudy (anonymous):

Use the formula: \[m=\frac{ y _{2}-y_{1} }{ x_{2}-x_{1}}\]

OpenStudy (anonymous):

ok brb imma write that down and work on it

OpenStudy (anonymous):

ok. Plug in the values for each point

OpenStudy (anonymous):

so is y2 = -4? and y1 =8?

OpenStudy (anonymous):

no, each point has an x and a y

OpenStudy (anonymous):

???

OpenStudy (anonymous):



OpenStudy (anonymous):

oooooohhhhhh ok hold up i think i got this...

OpenStudy (anonymous):

give me just a sec

OpenStudy (anonymous):

so m=-6-8/-3-(-4)?

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

the / is the middle line thingy

OpenStudy (anonymous):

division bar yes

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

what answer do you get?

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

2/1?

OpenStudy (anonymous):

Check your answer again

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2/7?

OpenStudy (anonymous):

is it the 2?

OpenStudy (anonymous):

No, you might want to rewrite the expression, since your parnthese are probably messing up your order of operations

OpenStudy (anonymous):

also make sure all of your signs are correct

OpenStudy (anonymous):

well ik that when u have a pos. and a neg. on bottom that u change the neg. and the minus sign on bottom to pos. # and addition sign

OpenStudy (anonymous):

yes, retry the problem and be careful to write it properly

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so m=-6-8/-3+4?

OpenStudy (anonymous):

that looks write but make sure to use parentheses to do the correct operations first

OpenStudy (anonymous):

so im doing -6-8 is 2 and -3+4 is 1 thats what im getting and since its wrong what is the wrong step that im making...?

OpenStudy (anonymous):

is -6-8 =2?

OpenStudy (anonymous):

is it the -6-8? yea i think its 48?

OpenStudy (anonymous):

please take another look. -6-8 is not 2

OpenStudy (anonymous):

-6+(-8)=?

OpenStudy (anonymous):

.... ok for some reason im lost i still have yet to understand this... :(

OpenStudy (anonymous):

ive been doing the lesson for this and just dont get it

OpenStudy (anonymous):

this is just arithmetic, slow down and look at what your missing.

OpenStudy (anonymous):

check your signs and it will come to you

OpenStudy (anonymous):

-6 + (-8) = ?

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

when you have two numbers with the same sign what do you do?

OpenStudy (anonymous):

either +14 or -14?

OpenStudy (anonymous):

There we go which is it?

OpenStudy (anonymous):

ik its one of them now

OpenStudy (anonymous):

how do you know this?

OpenStudy (anonymous):

um i did this part in 7th grade i think -14 but my mind wants me to say 14...

OpenStudy (anonymous):

why do you want 14 so badly?

OpenStudy (anonymous):

uugghh im never gonna get this...

OpenStudy (anonymous):

look at the signs and it will tell you.

OpenStudy (anonymous):

don't get frustrated. I promise the math will not lie to you

OpenStudy (anonymous):

idk maybe cuz im thinking * right now and when u have two negs or a pos and neg u make positive

OpenStudy (anonymous):

ah so that is only for multiplication and division

OpenStudy (anonymous):

so its -14?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

i knew it.. SHUT UP MIND AND LET ME DO MY WORK!!! lol

OpenStudy (anonymous):

hehe. well done

OpenStudy (anonymous):

ok so its -14/1?

OpenStudy (anonymous):

hey holly00d can u help me finish?

OpenStudy (anonymous):

nvmd